The average of the salaries of Kelly, Renee, and Nina is $38,000 a year. If Kelly earns $6000 less than Renee, and Nina's salary is two-thirds of Renee's salary, find the salary of each person
Answer by amarjeeth123(570)

The average of the salaries of Kelly, Renee, and Nina is $38,000 a year.
The total of the salaries of Kelly, Renee, and Nina is $114000 a year.
Let Renee's salary be x.
Kelly's salary is (x-6000).
Nina's salary is 2x/3.
We have x+(x-6000)+2x/3=114000
Simplifying we get
2x+2x/3=120000
8x/3=120000
x=120000*(3/8)=45000
Renee's salary is $45000.
Kelly's salary is $39000.
Nina's salary is $30000.